Using the LM-IS equilibrium, Explain how a change in a monetary variable could lead to a change in

the Real GDP?
Answer #1

In the above diagram we have output Y on X-axis and interest rate i on Y-axis. Then we have downward sloping IS curve and upward sloping LM curve.

As you might know IS curve is all combination of all the different interest rates and output level at which goods market is in equilibrium and LM curve is the combination of different interest rate and output level for the money market is in equilibrium.

IS Equation : Y = C + I + G +NX

Here C = private consumption

I = Investment

G = government expenditure

NX = net exports.

LM equation : L(Y, I) = M/P

Here L(Y, i) = demand function for real money balance

M/P = supply of real money balance.

Initially the economy is in equilibrium which is at point A with equilibrium output level of Y and equilibrium interest rate of i.

Now let's consider a change in monetary variable via monetary policy. Let's consider an expansionary monetary policy in which central bank increase the supply of nominal money, M.

Short run effect of increase in nominal money supply.

Let the original nominal money supply be M and original price level be P. Now central bank increases the nominal money supply to M' and in short run prices don't change so price level remains at P. So the new real money balance increase to M'/P.

This change is depicted by the rightwards shift in LM curve from LM to LM'. And since there is no change is fiscal variable the IS curve remains the same. With increased money supply the interest rate must fall in order to induce people to hold more money and decreased interest rate increases the investment.

The output increases and interest rate falls. The new short run equilibrium is restored at point B where both money and goods market is in equilibrium.

In short run real GDP increases because in short run there is no time for prices to adjust and hence only output increases but not prices. And as a result real GDP will increase in short run.

​​​​​​Effect of increased money supply in medium run.

In medium term prices are allowed to increase because there is enough time for all the economic agent to make changes in their behavior. As increased money supply lowers interest rate, the demand for investment spending goes up on demand side. Now on the supply side the supply must increase. Because initially the economy was in full employment level of equilibrium which means all those who wants to work at given wage rate can get work. But now the supply can only be increased by highring more labors and to do that firms must pay higher wages.

As firms increase the wages the cost of production goes up as result firms starts increasing the price of their goods and services. As result the general price level in economy in medium run increases. The increases in general price level, P will be exactly equal to what was increase in nominal money Supply, M. As a result the real money balance (M/P) decreases on account of increase in P.

In medium run, the real money balance reaches the previous value. Such that LM curves return to original position.

= M/P = M'/P'

The new LM curve equation becomes

= L(Y, i) = M'/P'

The result of all this is that LM curves shifts back to its original position and the economy reaches to the previous equilibrium point A. Where the equilibrium output is at Y and interest rate is at i.

But now the prices have increased from previous level. And since now the output is at previous level and prices have increased as a result the real GDP falls.

Originally, before increase in money supply output was at Y and price level was at P. The real GDP was Y/P.

And now the output level is Y same as previous and price level is at P' > P. The real GDP will be now Y/P'.

In medium run the real GDP falls.